# Vision-QRWKV: Exploring Quantum-Enhanced RWKV Models for Image Classification

### v1

@@ 4.1 Datasets
  • MedMNIST Collection [medmnistv2]: A curated benchmark for lightweight biomedical image classification, comprising 12 datasets covering modalities such as histopathology, chest X-ray, dermoscopy, breast ultrasound, and retina scans. We select the following: BloodMNIST, TissueMNIST, OCTMNIST, PathMNIST, ChestMNIST, OrganAMNIST, OrganSMNIST, OrganCMNIST, DermaMNIST, PneumoniaMNIST, RetinaMNIST, and BreastMNIST.
All images are resized to 28$\times$28 resolution to maintain consistency across datasets and enable direct comparison with MNIST-scale models.
@@ 5.1 Overall Performance Comparison
MNIST 96.49 96.12 No
BloodMNIST 91.32 92.22 Yes
FashionMNIST 85.56 86.08 Yes
PneumoniaMNIST 83.81 84.78 Yes
BreastMNIST 77.56 77.56 Equal
ChestMNIST 74.44 77.26 Yes
OrganAMNIST 78.24 76.27 No
OrganCMNIST 75.60 74.66 No
DermaMNIST 71.97 71.97 Equal
PathMNIST 74.00 71.11 No
OrganSMNIST 60.45 59.18 No
OCTMNIST 55.80 57.00 Yes
TissueMNIST 55.43 55.48 Yes
RetinaMNIST 49.25 53.75 Yes

### v2

@@ 4.1 Datasets
  • MedMNIST Collection [10]: A curated benchmark for lightweight biomedical image classification, comprising 12 datasets covering modalities such as histopathology, chest X-ray, dermoscopy, breast ultrasound, and retina scans. We select the following: BloodMNIST, TissueMNIST, OCTMNIST, PathMNIST, ChestMNIST, OrganAMNIST, OrganSMNIST, OrganCMNIST, DermaMNIST, PneumoniaMNIST, RetinaMNIST, and BreastMNIST.
All images are resized to 8$\times$8 resolution to maintain consistency across datasets and enable direct comparison with MNIST-scale models.
@@ 5.1 Overall Performance Comparison
MNIST [11] 96.49 96.12 No
BloodMNIST [10] 91.32 92.22 Yes
FashionMNIST [12] 85.56 86.08 Yes
PneumoniaMNIST [10] 83.81 84.78 Yes
BreastMNIST [10] 77.56 77.56 Equal
ChestMNIST [10] 74.44 77.26 Yes
OrganAMNIST [10] 78.24 76.27 No
OrganCMNIST [10] 75.60 74.66 No
DermaMNIST [10] 71.97 71.97 Equal
PathMNIST [10] 74.00 71.11 No
OrganSMNIST [10] 60.45 59.18 No
OCTMNIST [10] 55.80 57.00 Yes
TissueMNIST [10] 55.43 55.48 Yes
RetinaMNIST [10] 49.25 53.75 Yes